OpenStudy (wcrmelissa2001):

Help please! Question with inequalities (I've tried cases and squaring both sides but it doesn't work)

OpenStudy (wcrmelissa2001):

(1/(|x|-1))<=2/|x|

OpenStudy (prizzyjade):

\[1\div \left( \left| x\right| -1\right)\le 2\div \left| x \right|\]

OpenStudy (prizzyjade):

let right and left be equal \[\frac{ 1 }{ \left| x \right|-1 } = \frac{ 2 }{ x}\] cross multiply \[\left| x \right| = 2\left| x \right|-1\] transpose \[\left| x \right|=1\] then substitute the value of x by 1 \[0\le2\]

OpenStudy (phi):

did you move everything to the left side, and combine the fractions ?

OpenStudy (prizzyjade):

cross multiply

OpenStudy (phi):

you should get \[ \frac{2-|x| }{|x|(|x|-1)}\le 0 \] |x| is always positive. we need to figure out the signs for the other two factors.

OpenStudy (phi):

for the fraction to be negative, we require either 1) 2-|x| positive and |x|-1 negative or 2) 2-|x| negative and |x|-1 positive case 1) 2-|x| >= 0 |x|-1<0 (not =0 , because we do not want divide by 0) the first line gives: 2 >= |x| or |x| <= 2 second line gives: |x| < 1 for both to be true, we must have |x| < 1 this means: x <1 or -x < 1 --> x > -1 thus : -1 < x < 1 (note x \(\ne\) 0 because no divide by 0 allowed) case 2) 2-|x| <= 0 ---> |x| >= 2 |x| -1 > 0 ----> |x| >1 for both to be true at the same time, we use |x| >= 2 x <= -2 or x>= 2

OpenStudy (phi):

on a number line